Hi
The table below shows the number of cookies that were eaten from Monday to Friday. On Monday 2/3 as many cookies were eaten as on Friday. An average of 6 cookies per day were eaten. How many were eaten on Friday.
Monday x cookies
Tuesday 8 cookies
Wed 3 cookies
Thurs 4 cookies
Friday x cookies
Thank you

Answer by ewatrrr(24785)


Hi
Question States:
 ((2/3)x + 8 + 3 + 4 + x)/5 = 6
 ((2/3)x + 8 + 3 + 4 + x) = 30
  ((2/3)x  + x) = 15
      (5/3)x = 15
           x = 15(3/5) = 9, amount of cookies eaten on Friday

( 6 + 8 + 3 + 4 + 9)/5 = 6  checks
Wish You the Best in your Studies.